Evaluation of a polynomial in one variable: Evaluate the expression
b2 - 6b - 2
when b= -5
: Evaluation of a polynomial in one variable: Evaluate the expression
b2 - 6b - 2
when b= -5

Answer by jim_thompson5910(9897)

When nature calls, don't put her on speaker



b^2 - 6b - 2 Start with the given expression.


(-5)^2 - 6(-5) - 2 Plug in b=-5.


25-6(-5)-2 Square -5 to get 25.


25+30-2 Multiply -6 and -5 to get 30.


53 Combine like terms.


So (-5)^2 - 6(-5) - 2=53 which means that b^2 - 6b - 2=53 when b=-5